let V={(a,b,c)€R³|a+b=c} and W={(a,b,c)€R³|a=b} be subspaces of R³.Is R³ direct sum of V and W?


1
Expert's answer
2021-06-10T05:30:24-0400

Let "V=\{(a,b,c)\in\R^3\ |\ a+b=c\}" and "W=\{(a,b,c)\in\R^3\ |\ a=b\}" be subspaces of "\R^3".


Since "V=\{(a,b,c)\in\R^3\ |\ a+b=c\}=\{(a,b,a+b)\in\R^3\ |\ a,b\in\R\}=\{a(1,0,1)+b(0,1,1)\in\R^3\ |\ a,b\in\R\},"

we conclude that the vector subspace "V" is of dimension 2 with basis consisting of "(1,0,1)" and "(0,1,1)."


Taking into account that "W=\{(a,b,c)\in\R^3\ |\ a=b\}=\{(a,a,c)\in\R^3\ |\ a,c\in\R\}=\{a(1,1,0)+c(0,0,1)\in\R^3\ |\ a,c\in\R\},"

we conclude that the vector subspace "W" is of dimension 2 with basis consisting of "(1,1,0)" and "(0,0,1)."


Since "\dim V+\dim W=2+2=4>3=\dim \R^3," we conclude that "\R^3" can not be a direct sum of "V" and "W."